OpenStudy (anonymous):

simplifyy 9y^2-1/2y-1 X[multiply] 1-2y/3y-1

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

\[\frac{9x^2-1}{2y-1}\times \frac{1-2y}{3y-1}\]

OpenStudy (anonymous):

\[\frac{(3y-1)(3y+1)}{2y-1}\times \frac{1-2y}{3y-1}\]

OpenStudy (anonymous):

3y-1 cancels top and bottom and \[\frac{1-2y}{2y-1}=-1\] so it cancels as well

OpenStudy (anonymous):

leaving nothing but \[-(3y+1)=-3y-1\]

OpenStudy (anonymous):

so that means y is 1 ?

OpenStudy (anonymous):

thanks a lot